# Computational Hypergraph Discovery, a Gaussian Process framework for connecting the dots

### v1

@@ 6.2. How do we select the prior on the level of noise?
It follows that if the histogram of the eigenvalues of $D_{\gamma}$ is concentrated near 00 or near $1$, then the noise-to-signal ratio is non-informative since the prior $\gamma$ dominates it. To avoid this phenomenon, we select $\gamma$ so that the eigenvalues of $D_{\gamma}$ are well spread out in the sense that the histogram of its eigenvalues has maximum or near-maximum variance (see Fig. 19 for a good choice and a bad choice for $\gamma$). If the eigenvalues have an algebraic decay, then this is equivalent to taking $\gamma$ to be the geometric mean of those eigenvalues.
@@ 6.2.3. Rationale for the choices of $\gamma$
The purpose of this section is to present a rationale for the proposed choices for $\gamma$ in Sec. 6.2.1 and 6.2.2. For the choice Sec. 6.2.1, we present an asymptotic analysis of the signal-to-noise ratio in the setting of a simple linear regression problem. According to (6.8), $\gamma$ must scale linearly in $N$; this scaling is necessary to achieve a ratio that represents the signal-to-noise per sample. Without it (if $\gamma$ remains bounded as a function of $N$), this scaling of the signal-to-noise would converge towards 00 as $N\rightarrow\infty$. To see how we will now consider a simple example in which we seek to linearly regress the variable $y$ as a function of the variable $x$, both taken to be scalar (in which case $\psi(x)=x$). Assume that the samples are of the form $Y_{i}=aX_{i}+\sigma Z_{i}$ for $i=1,\ldots,N$, where $a,\sigma\not=0$, the $Z_{i}$ are i.i.d. $\mathcal{N}(0,1)$ random variables, and the $X_{i}$ satisfy $\frac{1}{N}\sum_{i=1}^{N}X_{i}=0$ and $\frac{1}{N}\sum_{i=1}^{N}X_{i}^{2}=1$. Then, the signal-to-noise ratio is $\frac{\mathcal{V}(s)}{\mathcal{V}(s)+\mathcal{V}(n)}$ with $\mathcal{V}(s)=|v|^{2}$ and $\mathcal{V}(n)=\frac{1}{\gamma}\sum_{i=1}^{N}|vX_{i}-Y_{i}|^{2}$ and $v$ is a minimizer of
@@
which converges towards 00 as $\sigma\rightarrow\infty$ and towards $1/(1+\sigma^{2})$ as $a\rightarrow\infty$, which has, therefore, the desired properties.
@@
If the noise is only comprised of noise, then an interval of confidence can be obtained on the noise-to-signal ratio. To describe this consider the problem of testing the null hypothesis ${\bf H_{0}}:f^{\dagger}\equiv 0$ (there is no signal) against the alternative hypothesis ${\bf H_{1}}:f^{\dagger}\not\equiv 0$ (there is a signal). Under the null hypothesis ${\bf H_{0}}$, the distribution of the noise-to-signal ratio (6.9) is known and it follows that of the random variable
@@
7. Numerical experiments.
We will now illustrate the proposed approach through numerical experiments that can be reproduced in the Github repository of the paper.
Figure 20. Algebraic equations examples.
@@
The examples discussed in this section (Fig. 20.(a-d)) illustrate the application of the proposed approach to the recovery of functional dependencies from data satisfying hidden algebraic equations. In all these examples, we have $d=6$ or $d=7$ variables and $N=1000$ samples from those variables. For $d=6$ the variables are $w_{1},w_{2},w_{3},w_{4},x_{1},x_{2}$. For $d=7$ the variables are $w_{1},w_{2},w_{3},w_{4},x_{1},x_{2},x_{3}$. The samples from the variables $w_{1}$ to $w_{4}$ are i.i.d. $\mathcal{N}(0,1)$ random variables, and the samples from $x_{1}$, $x_{2}$ (and $x_{3}$ for $d=7$) are functionally dependent on the other variables. Although we have used Alg. 2 for the algebraic equations examples presented in this section, Alg. 1 yields the same results with the default signal-to-noise threshold $\tau=0.5$.
7.1.1. Fig. 20.(a)
In this example, $d=6$ and the samples from $x_{1}$ and $x_{2}$ satisfy the equations
$x_{1}=w_{1}\text{ and }x_{2}=w_{2}\,.$ (7.1)
Fig. 20.(a) shows the hypergraph recovered by the proposed approach using Alg. 2. The recovery is accurate and illustrates the one-to-one mappings between $x_{1}$ and $w_{1}$, as well as between $x_{2}$ and $w_{2}$.
7.1.2. Fig. 20.(b)
In this example, $d=7$ and the samples from $x_{1},x_{2}$ and $x_{3}$ satisfy the equations
$x_{1}=w_{1},\quad x_{2}=x_{1}^{2}+1+0.1w_{2},\text{ and }x_{3}=w_{3}\,.$ (7.2)
Fig. 20.(b) shows the hypergraph recovered by the proposed approach using Alg. 2. The results are notably accurate. The algorithm rightly pinpoints the direct correspondence between $x_{1}$ and $w_{1}$, as well as between $x_{3}$ and $w_{3}$. Yet, the efficiency of the algorithm is compromised when duplicate variables are present. For instance, given the quadratic kernel, even though $x_{2}$ can trace back its origin to either $x_{1}$ and $w_{2}$ or $w_{1}$ and $w_{2}$, the algorithm recognizes $x_{1}$, $w_{1}$, and $w_{2}$ as its ancestors. This underscores the significance of eliminating redundant variables when aiming to derive the sparsest and the most understandable graph.
7.1.3. Fig. 20.(c)
In this example, $d=6$ and the samples from $x_{1}$ and $x_{2}$ satisfy the equations
$x_{1}=w_{1}w_{2}\text{ and }x_{2}=w_{2}\sin(w_{4})\,.$ (7.3)
Fig. 20.(c) shows the hypergraph recovered by the proposed approach using Alg. 2. The recovery is accurate. The underlying functional dependencies are inherently nonlinear and were not explicitly represented in our feature maps. By utilizing the fully nonlinear kernel (4.31), our approach identified the correct functional dependencies.
7.1.4. Fig. 20.(d)
In this example, $d=7$ and the samples from $x_{1},x_{2}$ and $x_{3}$ satisfy the equations
$x_{1}=w_{1},\quad x_{2}=x_{1}^{3}+1+0.1w_{2}\text{ and }x_{3}=(x_{1}+2)^{3}+0.1w_{3}\,.$ (7.4)
Fig. 20.(d) shows the hypergraph recovered with Alg. 2. Given our kernel selections (linear, quadratic, and fully nonlinear), one might anticipate an approximate recovery of the equations via the fully nonlinear kernel, especially considering the $x_{1}^{3}$ term. However, the method astutely identifies an alternative combination, permitting an exact recovery with the quadratic kernel. Notably, upon expanding $(x_{1}+2)^{3}$, it becomes evident that the cubic terms can be portrayed using linear and quadratic combinations of other nodes. Such an automatic simplification in equations is undeniably profound. Such automatic discovery of simplifying equations is very powerful. Yet, it is worth mentioning that to achieve this exact quadratic recovery; we must consider more ancestors than originally denoted in (7.4). This emphasizes the notion that functional dependencies are profoundly influenced by the chosen function form, and our methodology discerns one of the potential dependencies.
Figure 21. Output graph of the chemistry example. We observe perfect recovery of the equations in (7.5), which are also correctly identified as quadratic.
7.2. Chemical reaction network.
We now consider the example described in Sec. 1.3.1 and taken from [19]. The proposed mechanism for the hydrogenation of ethylene $(\text{C}_{2}\text{H}_{4})$ to ethane $(\text{C}_{2}\text{H}_{6})$, is (writing $[H]$ for the concentration of $H$) modeled by the following system of differential equations
@@
The primary variables are the concentrations $[H_{2}]$, $[H]$, $[C_{2}H_{4}]$ and $[C_{2}H_{5}]$ and their time derivatives $\frac{d[H_{2}]}{dt}$, $\frac{d[H]}{dt}$, $\frac{d[C_{2}H_{4}]}{dt}$ and $\frac{d[C_{2}H_{5}]}{dt}$. The computational hypergraph encodes the functional dependencies (7.5) associated with the chemical reactions. The hyperedges of the hypergraph are assumed to be unknown and the primary variables are assumed to be known. Given $N$ samples from the graph of the form
@@
our objective is to recover the structure of the hypergraph given by (7.5), representing the functions by hyperedges. We create a dataset of the form (7.6) by integrating 50 trajectories of (7.5) for different initial conditions, and each equispaced 50 times from $t=0$ to $t=5$. The dataset is represented in Figure 5 (the time derivatives of concentrations are estimated by taking the derivatives of the interpolants of those concentrations). We impose the information that the derivative variables are function of the non-derivative variables to avoid ambiguity in the recovery, as (7.5) is not the unique representation of the functional relation between nodes in the graph. We implement Alg. 1 with weights $\beta=[0.1,0.01,0.001]$ for linear, quadratic, and nonlinear, respectively (Alg. 2 recovers the same hypergraph). The output graph can be seen in Fig. 21. We obtain a perfect recovery of the computational graph and a correct identification of the relations being quadratic.
Figure 22. (a) Full retrieved graph. (b) Policy subgraph (c) “New intensive care patients” subgraph. (d) Recovered graph from the COVID-19 open data after removing redundant information.
@@
This section uses the CHD framework as a statistical analysis tool to scrutinize COVID-19 open data, as detailed in Sec. 1.3.2. Categorical data are treated as scalar values, with all variables scaled to achieve a mean of 0 and a variance of 1. We implement three distinct kernel types: linear, quadratic, and Gaussian, with a length scale of 1 for the latter. A weight ratio of $1/10$ is assigned between kernels, signifying that the quadratic kernel is weighted ten times less than the linear kernel. Lastly, the noise parameter, $\gamma$, is determined using the optimal value outlined in Sec. 6.
Initially, a complete graph is constructed using all variables, depicted in Fig.22.(a). This construction is done using only linear and quadratic kernels. This is done to regularise our problem by choosing a feature map whose dimension is not too high. An example of the evolution of the signal-to-noise ratio we observe is given in Fig. 23. In this case, it is very clear where we need to stop pruning, as the jumping noise ratio is significant. A notably clustered section within the graph is observed upon examination. Fig.22.(b) offers an in-depth view of this cluster, revealing its representation of the government’s pandemic response. This evident clustering leads to the conclusion that the government either implemented multiple restrictive measures simultaneously or lifted them in unison. From the subgraph shown in Fig.22.(b), a few representative nodes for the cluster are selected, ones with few and serving as an ancestor to numerous nodes. The vaccination policy and stay-at-home requirements are deemed an apt choice in this context, especially given their impact on the population. Additionally, new vaccinations form a cluster (Fig. 22.(c)), displaying a linear relationship between nodes, signaling redundant information. Eliminating redundant nodes is vital for two reasons: firstly, it improves the graph’s readability, especially with 31 variables; secondly, it avoids hindering graph discovery. In an extreme case, treating two identical variables as distinct would result in one variable’s ancestor simply being its duplicate, yielding an uninformative graph.
Figure 23. Evolution of the noise-to-signal ratio when pruning ancestors for the cumulative number of hospitalized patients.
Subsequently, the graph discovery algorithm is rerun, with reduced variables due to eliminating redundancy, ushering us into a predominantly noisy regime. With fewer variables available, we use additionally the Gaussian kernel. Two indicators are employed to navigate our discovery process: the signal-to-noise ratio and the Z-test. The former quantifies the degree to which our regression is influenced by noise, while the latter signals the existence of any signal. We follow the procedure in algorithm 2, resulting in the graph presented in Fig. 22.(e). Here, some ancestor relationships are logical when interpreting using our experience of the COVID-19 pandemic. For instance, the number of new hospitalized and new deceased appear closely linked. Indeed, most recorded deaths are from people in a hospital. Additionally, it is observed that many ancestors are needed to explain the number of people tested. This number depends not only on the state of the epidemic but also on the government’s and the population’s reaction. Furthermore, many tests are conducted during holidays such as Christmas, regardless of the current pandemic status. This widespread testing complicates the ability to predict testing based solely on present variables. However, it is vital to acknowledge that not all relationships in the graph necessarily suggest causal connections. Our method determines graph edges based on functional relationships, not causal ones. Ultimately, this graph demonstrates sparsity, facilitating a superior understanding of the data’s underlying structure. Furthermore, the method allows us to identify and eliminate redundant information, enhancing clarity regarding the most pivotal trends observed.
@@ 7.4. Cell signaling network
Lastly, we apply the CHD framework to discover a hierarchy of functional dependencies in biological cellular signaling networks. We consider the problem discussed in Sec. 1.3.4 and illustrated in Fig. 7.(b). In this experiment, we use single-cell data consisting of the $d=11$ phosphoproteins and phospholipids levels that were measured using flow cytometry. This dataset was studied from a probabilistic modeling perspective in previous works. While [34] learned a directed acyclic graph to encode causal dependencies, [13] learned an undirected graph of conditional independencies between the $d$ molecule levels by assuming the underlying data follows a multivariate Gaussian distribution. The latter analysis encodes acyclic dependencies but does not identify directions. In this work, we aim to identify the functional dependencies without imposing strong distributional assumptions on the data.
To identify the ancestors of each node, we apply the algorithm in two stages. First, we learn the dependencies using only linear and quadratic kernels. Figure 24 left identifies the resulting graph learned given a subset of $N=2,000$ samples chosen uniformly at random from the dataset. We observe that the graph identified by the algorithm consists of four disconnected clusters where the molecule levels in each cluster are closely related by linear or quadratic dependencies (all connections are linear except for the connection between Akt and PKA, which is quadratic). These edges match a subset of the edges found in the gold standard model identified in [34]. With perfect dependencies that have no noise, one can define constraints that reduce the total number of variables in the system. For this noisy dataset that, we treat these dependencies as forming groups of similar variables and introduce a hierarchical approach to learn the connections between groups.
Figure 24. (Left) Graph identified based on linear and quadratic kernels. (Right) Graph identified with the proposed multi-level approach. Solid arrows represent the strong intra-cluster connections identified in the first level, and the dashed lines represent the relatively weaker connections between nodes and clusters identified in the second level. The width and grayscale intensities of each edge correspond to its signal-to-noise ratio.
Figure 25. (Left) Graph identified in [34] where the nodes are colored based on their membership in a cluster. (Right) Graph identified with the proposed multi-level approach showing similar intra-cluster connections as the left figure as well as the dominant connections outside each cluster. We remind the reader that the arrows identified by the CHD framework encode functional dependence rather than causality.
@@
Second, we run the graph discovery algorithm after grouping the molecules into clusters. For each node in the graph, we identified the ancestors of each node by constraining the dependence to be a subset of the clusters. In other words, when identifying the ancestors of a given node $i$ in cluster $C$, the algorithm is only permitted to (1) use ancestors that do not belong to cluster $C$, and (2) include all or none of the variables in each cluster ($j$ in cluster $D\not=C$ is listed as an ancestor if and only if all other nodes $j^{\prime}$ in cluster $D$ are also listed as ancestors). The ancestors were identified using a Gaussian (fully nonlinear) kernel and the number of by ancestors were selected manually based on the inflection point in the noise-to-signal ratio. The resulting graph is depicted on the right side of Figure 24. Each edge is weighted based on its signal-to-noise ratio. We observe that there is a stronger dependence of the Jnk, PKC, and P38 cluster on the PIP3, Plcg, and PIP2 cluster, which closely matches the gold standard model. As compared to approaches based on acylic DAGs, however, the graph identified by our algorithm also contains feedback loops between the various molecule levels. Fig. 25 displays a side-by-side comparison between the graph generated in [34] and the graph identified with our method. To aid in this comparison, we have highlighted different clusters in distinct colors. We emphasize that while the Bayesian network analysis in [34] relied on the control of the sampling of the underlying variables (the simultaneous measurement of multiple phosphorylated protein and phospholipid components in thousands of individual primary human immune system cells, and perturbing these cells with molecular interventions), the reconstruction obtained by our method did not use this information and recovered functional dependencies rather than causal dependencies. Interestingly, the information recovered through our method appears to complement and enhance the findings presented in  [34] (e.g., the linear and noiseless dependencies between variables in the JNK cluster is not something that could easily be inferred from the graph produced in [34]).
7.5. Discussions
In our numerical experiments, we observe that the proposed algorithm maintains robustness with respect to the selection of the signal kernel, i.e., the $\beta_{i}$ in the kernel (2.7). Furthermore, the noise prior variance $\gamma$ identified by our algorithm is very close to the regularization term that minimizes cross-validation mean square loss. More precisely, the identification of functional dependencies between variables relies on regressing input-output data $(X,Y)$ for the signal function $f$ with a kernel $K$ and a nugget term $\gamma$ ($f(\cdot)=K(\cdot,X)\big{(}K(X,X)+\gamma I\big{)}^{-1}Y$). Writing $(X^{c},Y^{c})$ for a subset of the input-output data, $f^{c}(\cdot)=K(\cdot,X^{c})\big{(}K(X^{c},X^{c})+\gamma I\big{)}^{-1}Y^{c}$ for the corresponding kernel ridge regressor, and $\gamma_{c}$ for the minimizer of $L^{2}$ error $\|Y-f^{c}(X)\|$, the value of $\gamma$ selected by our algorithm is very close to $\gamma_{c}$. This suggests that one could also use cross-validation (applied to recover the signal function between a given node and a group of nodes in computing a noise-to-signal ratio) to select $\gamma$.
We also observe that forming clusters from highly interdependent variables helps to obtain a sparser graph. Additionally, the precision of the pruning process is enhanced by avoiding the division of node activation within the cluster among its separate constituents. We employed this strategy in the recovery of the gene expression graph in Sec. 7.4. See Sec. 7.1.2 for an algebraic equation example where this strategy would have enhanced the recovery but was not used. We also note from the algebraic equation recovery examples of Sec. 7.1.4 that we can only hope to recover equivalent equations, and the type of equations recovered depends on the kernel employed for the signal.
8. Conclusion
We have developed a comprehensive Gaussian Process framework for solving Type 3 (hypergraph discovery) problems, which is interpretable and amenable to analysis. The breadth and complexity of Type 3 problems significantly surpass those encountered in Type 2 (hypergraph completion), and the initial numerical examples we present serve as a motivation for the scope of Type 3 problems and the broader applications made possible by this approach. Our proposed algorithm is designed to be fully autonomous, yet it offers the flexibility for manual adjustments to refine the graph’s structure recovery, as detailed Sec. 7.4. Additionally, our method eliminates the need for a predetermined ordering of variables, a common requirement in acyclic probabilistic models where determining an optimal order is an NP-hard problem usually tackled using heuristic approaches. Furthermore, our approach can actually be utilized to generate such an ordering by quantifying the strength of the connections it recovers. The Uncertainty Quantification properties of the underlying Gaussian Processes are integral to the method and could also be employed to quantify uncertainties in the structure of the recovered graph. While the uncertainty quantification (UQ) properties of Gaussian Processes (GPs) lack a direct equivalent in Deep Learning, generalizing our approach to encompass richly hyperparametrized kernels could facilitate its extension to Deep Learning type connections between variables. This is achievable by interpreting neural networks as hyperparameterized feature maps.
Acknowledgements
HO, TB, PB, XY, and RB acknowledge support from the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation). Additionally, HO, TB, and PB acknowledge support by the Department of Energy under award number DE-SC0023163 (SEA-CROGS: Scalable, Efficient and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems) and by the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. HO and PB further acknowledge support by Beyond Limits (Learning Optimal Models) through CAST (The Caltech Center for Autonomous Systems and Technologies). TB acknowledges support from the Kortschak Scholar Fellowship Program. NR acknowledges support from the JPL Researchers on Campus (JROC) program.

### v2

@@ 7.2. How do we select the prior on the level of noise?
It follows that if the histogram of the eigenvalues of $D_{\gamma}$ is concentrated near 0 or near $1$, then the noise-to-signal ratio is non-informative since the prior $\gamma$ dominates it. To avoid this phenomenon, we select $\gamma$ so that the eigenvalues of $D_{\gamma}$ are well spread out in the sense that the histogram of its eigenvalues has maximum or near-maximum variance (see Fig. 6 for a good choice and a bad choice for $\gamma$). If the eigenvalues have an algebraic decay, then this is equivalent to taking $\gamma$ to be the geometric mean of those eigenvalues.
@@ 7.2.3. Rationale for the choices of $\gamma$
The purpose of this section is to present a rationale for the proposed choices for $\gamma$ in Sec. 7.2.1 and 7.2.2. For the choice Sec. 7.2.1, we present an asymptotic analysis of the signal-to-noise ratio in the setting of a simple linear regression problem. According to (7.8), $\gamma$ must scale linearly in $N$; this scaling is necessary to achieve a ratio that represents the signal-to-noise per sample. Without it (if $\gamma$ remains bounded as a function of $N$), this scaling of the signal-to-noise would converge towards 0 as $N\rightarrow\infty$. To see how we will now consider a simple example in which we seek to linearly regress the variable $y$ as a function of the variable $x$, both taken to be scalar (in which case $\psi(x)=x$). Assume that the samples are of the form $Y_{i}=aX_{i}+\sigma Z_{i}$ for $i=1,\ldots,N$, where $a,\sigma\not=0$, the $Z_{i}$ are i.i.d. $\mathcal{N}(0,1)$ random variables, and the $X_{i}$ satisfy $\frac{1}{N}\sum_{i=1}^{N}X_{i}=0$ and $\frac{1}{N}\sum_{i=1}^{N}X_{i}^{2}=1$. Then, the signal-to-noise ratio is $\frac{\mathcal{V}(s)}{\mathcal{V}(s)+\mathcal{V}(n)}$ with $\mathcal{V}(s)=|v|^{2}$ and $\mathcal{V}(n)=\frac{1}{\gamma}\sum_{i=1}^{N}|vX_{i}-Y_{i}|^{2}$ and $v$ is a minimizer of
@@
which converges towards 0 as $\sigma\rightarrow\infty$ and towards $1/(1+\sigma^{2})$ as $a\rightarrow\infty$, which has, therefore, the desired properties.
@@
If the data is only comprised of noise, then an interval of confidence can be obtained on the noise-to-signal ratio. To describe this consider the problem of testing the null hypothesis ${\bf H_{0}}:f^{\dagger}\equiv 0$ (there is no signal) against the alternative hypothesis ${\bf H_{1}}:f^{\dagger}\not\equiv 0$ (there is a signal). Under the null hypothesis $\bf H_{0}$}, the distribution of the noise-to-signal ratio (7.9) is known and it follows that of the random variable
@@
8. Supplementary information on examples.
@@
Although we have used Alg. 2 for the algebraic equations examples presented in Fig. 4, Alg. 1 yields the same results with the default signal-to-noise threshold $\tau=0.5$.
8.2. The chemical reaction network.
Consider the chemical reaction network example illustrated in Fig. 4.(a). The proposed mechanism for the hydrogenation of ethylene $(\text{C}_{2}\text{H}_{4})$ to ethane $(\text{C}_{2}\text{H}_{6})$, is (writing $[H]$ for the concentration of $H$) modeled by the following system of differential equations
@@
The primary variables are the concentrations $_{2}], [H][H$], $C_{2}H_{4}]$ and $C_{2}H_{5}]$ and their time derivatives $\frac{d[H_{2}]}{dt}$, $\frac{d[H]}{dt}$, $\frac{d[C_{2}H_{4}]}{dt}$ and $\frac{d[C_{2}H_{5}]}{dt}$. The computational hypergraph encodes the functional dependencies (8.1) associated with the chemical reactions. The hyperedges of the hypergraph are assumed to be unknown and the primary variables are assumed to be known. Given $N$ samples from the graph of the form
@@
our objective is to recover the structure of the hypergraph given by (8.1), representing the functions by hyperedges. We create a dataset of the form (8.2) by integrating 50 trajectories of (8.1) for different initial conditions, and each equispaced 50 times from $t=0$ to $t=5$. The dataset is represented in Fig. 4.(b) (the time derivatives of concentrations are estimated by taking the derivatives of the interpolants of those concentrations). We impose the information that the derivative variables are function of the non-derivative variables to avoid ambiguity in the recovery, as (8.1) is not the unique representation of the functional relation between nodes in the graph. We implement Alg. 1 with weights $\beta=[0.1,0.01,0.001]$ for linear, quadratic, and nonlinear, respectively (Alg. 2 recovers the same hypergraph). The output graph can be seen in Fig. 4.(b). We obtain a perfect recovery of the computational graph and a correct identification of the relations being quadratic.
@@
Consider the example illustrated in Fig. 3.(e-k). Categorical data are treated as scalar values, with all variables scaled to achieve a mean of 0 and a variance of 1. We implement three distinct kernel types: linear, quadratic, and Gaussian, with a length scale of 1 for the latter. A weight ratio of $1/10$ is assigned between kernels, signifying that the quadratic kernel is weighted ten times less than the linear kernel. Lastly, the noise parameter, $\gamma$, is determined using the optimal value outlined in Sec. 7. Initially, a complete graph is constructed using all variables, depicted in Fig. 3.(g). This construction is done using only linear and quadratic kernels. The full graph is highly clustered and redundant information is eliminated by selecting representative nodes for each cluster. Eliminating redundant nodes is important for two reasons: firstly, it improves the graph’s readability, especially with 31 variables; secondly, it avoids hindering graph discovery. In an extreme case, treating two identical variables as distinct would result in one variable’s ancestor simply being its duplicate, yielding an uninformative graph. Subsequently, the graph discovery algorithm is rerun, with reduced variables due to eliminating redundancy, ushering us into a predominantly noisy regime. With fewer variables available, we use additionally the nonlinear kernel. Two indicators are employed to navigate our discovery process: the signal-to-noise ratio and the Z-test. The former quantifies the degree to which our regression is influenced by noise, while the latter signals the existence of any signal. We follow the procedure in algorithm 2, resulting in the graph presented in Fig. 3.(k).
@@ 8.4. Cell signaling network
Consider the example Fig. 1.(l) from [36] and Fig. 4.(h-j). To identify the ancestors of each node, we apply the algorithm in two stages. First, we learn the dependencies using only linear and quadratic kernels. Fig. 4.(h) identifies the resulting graph learned given a subset of $N=2,000$ samples chosen uniformly at random from the dataset. We observe that the graph identified by the algorithm consists of four disconnected clusters where the molecule levels in each cluster are closely related by linear or quadratic dependencies (all connections are linear except for the connection between Akt and PKA, which is quadratic). These edges match a subset of the edges found in the gold standard model identified in [36]. With perfect dependencies that have no noise, one can define constraints that reduce the total number of variables in the system. For this noisy dataset that, we treat these dependencies as forming groups of similar variables and introduce a hierarchical approach to learn the connections between groups. Second, we run the graph discovery algorithm after grouping the molecules into clusters. For each node in the graph, we identified the ancestors of each node by constraining the dependence to be a subset of the clusters. In other words, when identifying the ancestors of a given node $i$ in cluster $C$, the algorithm is only permitted to (1) use ancestors that do not belong to cluster $C$, and (2) include all or none of the variables in each cluster ($j$ in cluster $D\not=C$ is listed as an ancestor if and only if all other nodes $j^{\prime}$ in cluster $D$ are also listed as ancestors). The ancestors were identified using a Gaussian (fully nonlinear) kernel and the number of by ancestors were selected manually based on the inflection point in the noise-to-signal ratio. The resulting graph is depicted in Fig. 4.(i). Each edge is weighted based on its signal-to-noise ratio. We observe that there is a stronger dependence of the Jnk, PKC, and P38 cluster on the PIP3, Plcg, and PIP2 cluster, which closely matches the gold standard model. As compared to approaches based on acyclic DAGs, however, the graph identified by our algorithm also contains feedback loops between the various molecule levels. Fig. 4.(i-j) displays a side-by-side comparison between the graph identified with our method and the graph generated in [36]. To aid in this comparison, we have highlighted different clusters in distinct colors. We emphasize that while the Bayesian network analysis in [36] relied on the control of the sampling of the underlying variables (the simultaneous measurement of multiple phosphorylated protein and phospholipid components in thousands of individual primary human immune system cells, and perturbing these cells with molecular interventions), the reconstruction obtained by our method did not use this information and recovered functional dependencies rather than causal dependencies. Interestingly, the information recovered through our method appears to complement and enhance the findings presented in  [36] (e.g., the linear and noiseless dependencies between variables in the JNK cluster is not something that could easily be inferred from the graph produced in [36]).
8.5. BCR reaction network
In the high-dimensional example of the BCR reaction network, the computations of terms of the form $y^{T}k_{o}(X,X)y$ (i.e., the activations), where $y\in\mathbb{R}^{n}$ and $k_{o}(X,X)$ is the $o$-th coordinate of the quadratic kernel ($k(x_{i},x_{j})=(1+\langle x_{i},x_{j}\rangle)^{2}$) becomes the computational bottleneck of our method. If we let $x_{1},\ldots,x_{n}\in\mathbb{R}^{p}$ be the points and $x_{i}^{o}$ be the $o$-th coordinate of $x_{i}$, we can compute the activation of the $o$-th coordinate using
$k_{o}(x_{i},x_{j})=(1+x^{o}_{i}x^{o}_{j})^{2}-1+2x^{o}_{i}x^{o}_{j}\langle x_{i}^{-o},x_{j}^{-o}\rangle$ (8.3)
where $k_{o}$ is the $o$-th coordinate of the kernel and $x^{-o}_{i}$ represents the remaining coordinates of $x_{i}$. To compute the $n\times n$ kernel matrix of $k_{o}$ for each $o\in\{1,..,p\}$, we must compute $p\times n\times n$ inner products in $\mathbb{R}^{p}$, which is a very large computation. Instead, we may use the following reformulation to speed up computations. Notice $\langle x_{i},x_{j}\rangle=x_{i}^{o}x_{j}^{o}+\langle x_{i}^{-o},x_{j}^{-o}\rangle$, and therefore $k_{o}(x_{i},x_{j})=2x_{i}^{o}x_{j}^{o}\langle x_{i},x_{j}\rangle+2x_{i}^{o}x_{j}^{o}-(x_{i}^{o}x_{j}^{o})^{2}$. Now, define $v^{o}=(x_{i}^{o}y_{i})_{i=1}^{p}$ and $w^{o}=((x_{i}^{o})^{2}y_{i})_{i=1}^{p}$, and note that
$y^{T}K_{o}y=\sum_{i,j}2y_{i}x_{i}^{o}y_{j}x_{j}^{o}(1+\langle x_{i},x_{j}\rangle)-\sum_{i,j}y_{i}y_{j}(x_{i}^{o}x_{j}^{o})^{2}$ (8.4)
and so defining $\tilde{K}=(2(1+\langle x_{i},x_{j}\rangle))_{i,j=1}^{n}$ we have that
$y^{T}K_{o}y={v^{o}}^{T}\tilde{K}v^{o}-\left(\sum_{i=1}^{p}w_{i}^{o}\right)^{2}$ (8.5)
Note that $\tilde{K}$ is computed just once for all $p$, and only $v^{o}$ and $w^{o}$ change for every ancestor calculation, which is where the main computational gain comes from. One may find in the GitHub repository of the paper a comparison of the two methods of computations and observe a tenfold speedup. This speedup is even larger in our implementation of the BCR example, as GPU acceleration enables the second method to run even faster.